OpenStudy (anonymous):

Will give a medal: Determine whether the sequence converges or diverges. If it converges, give the limit. 108, -18, 3, negative one divided by two, ... Just need an explanation on how to do these types of problems.

OpenStudy (irishboy123):

Find the common ratio first

OpenStudy (anonymous):

@irishboy123 I didn't think there was one.

OpenStudy (anonymous):

@LynFran Could you please help?

OpenStudy (irishboy123):

\(108 \times ?? = -18\) \(-18 \times ?? = 3\) \( 3 \times ?? = -\frac{1}{2} \)

OpenStudy (anonymous):

-1/6 @IrishBoy123

OpenStudy (anonymous):

So then what? @IrishBoy123

OpenStudy (irishboy123):

yes so you took 108 and started multiplying it by -1/6 so the terms of the sequence gradually becomes what?

OpenStudy (anonymous):

0?

OpenStudy (irishboy123):

yes

OpenStudy (anonymous):

So converges at 0?

OpenStudy (anonymous):

Can you help me with one more? Write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 + ... + 67

OpenStudy (anonymous):

And it converges at 0 because the values get closer and closer to 0

OpenStudy (irishboy123):

yes for next one, post new thread

OpenStudy (anonymous):

Ok